OpenStudy (anonymous):

What is the axis of symmetry for the graph of y = 3x^2 + 6x − 4? x = 1 x = −1 x = 2 x = −2

OpenStudy (anonymous):

@Nurali

OpenStudy (nurali):

Compare with the general quadratic equation y = ax2 + bx + c. Here, a = –3, b = –6, and c = 4. Substitute a = –3 and b = –6 in the equation of axis of symmetry, x = – b/2a. x = –1 The equation of axis of symmetry is x = –1.